OpenStudy (anonymous):

how do you find the x-intercept or zero's of f(x)=(x-5)(2x+3)

OpenStudy (anonymous):

find for what values of x f(x)=0

OpenStudy (anonymous):

Set both "parenthesis" equal to zero and solve. So: x-5=0 and 2x+3=0. Solve for both and you have your answer

OpenStudy (anonymous):

then how do you find the y intercept for f(x)=2x-10

OpenStudy (anonymous):

believe it or not its actually -10. Why? this function is in the form of y=mx+b... where m is the slope and b is the y-intercept

OpenStudy (anonymous):

well how do you solve it?

OpenStudy (anonymous):

you dont. if you are just trying to find the y-intercept. then your answer is just -10. y-intercept is where the graph passes the y-axis. and it does so where y=-10

OpenStudy (anonymous):

oh ha ok thanks